OpenStudy (anonymous):

How to rewrite the this as Ax+By=c f(x)=3/2 (x-4)

OpenStudy (abmon98):

y=3/2(x-4) Get rid of the fraction by multiplying by 2 2y=3(x-4) 2y=3x-12 -3x+2y=-12

OpenStudy (anonymous):

can you show me the process of how did u get 2y=3(x-4)

OpenStudy (anonymous):

the numerator is 3 and the dinaminator is 2

OpenStudy (sheraz12345):

He just cross multiplied the terms @shiqki017

OpenStudy (anonymous):

can u show me that

OpenStudy (sheraz12345):

@Abmon98 how did u remove the fraction... it should be like this